Anna has 12 bills in her wallet, some $5 an some $10. The total value of the bills is $100.How many of each bill does Anna have?
Answer by scott8148(6628)

let x=tens, so 12-x=fives

10(x)+5(12-x)=100 ___ 10x+60-5x=100